OpenStudy (anonymous):

Worded, simultaneous equation plus ratios. I am looking to see whether this is an appropraite model, I don't really want the answer. If not, why not? Hard to get my head around... Okay. A pebble-mix company keeps three types, A, B and C, of pebble-mix in which the proportions of cement, pebbles and sand are respectively, 1:3:2, 1:4:6.5, 1:2:4. If an order is made by a builder for 26 tonnes of pebble-mis in which the ratio is 1:3.5:4.5, determine the amount of each pebble-mix. I tried to work it out with the following equations: Which I will writ in post as I am about to run out of space.

OpenStudy (anonymous):

1/6A + 1/11.5B + 1/7C = 26/9 3/6A + 4/11.5B + 2/7C = 91/9 2/6A + 6.5/11.5B + 4/7C = 117/9

OpenStudy (anonymous):

Also part of the trouble is I am unable to visualise it at all because I have no idea have to graph such a thing.

OpenStudy (anonymous):

I know how to graph 2 variables and I find it helps to understand those questions, but not 3 variables.

OpenStudy (anonymous):

I tried to work it out but I got the wrong answer, but I have no idea if the model is wrong or the working out I was trying.